# Initialization Schemes for Kolmogorov–Arnold Networks: An Empirical Study

### v2

@@ Abstract
Kolmogorov–Arnold Networks (KANs) are a recently introduced neural architecture that replace fixed nonlinearities with trainable activation functions, offering enhanced flexibility and interpretability. While KANs have been applied successfully across scientific and machine learning tasks, their initialization strategies remain largely unexplored. In this work, we study initialization schemes for spline-based KANs, proposing two theory-driven approaches inspired by LeCun and Glorot, as well as an empirical power-law family with tunable exponents. Our evaluation combines large-scale grid searches on function fitting and forward PDE benchmarks, an analysis of training dynamics through the lens of the Neural Tangent Kernel, and evaluations on a subset of the Feynman dataset. Our findings indicate that the Glorot-inspired initialization significantly outperforms the baseline in parameter-rich models, while power-law initialization achieves the strongest performance overall, both across tasks and for architectures of varying size. All code and data accompanying this manuscript are publicly available at https://github.com/srigas/KAN_Initialization_Schemes.
@@ Acknowledgments
S. R. and G. A. are supported by the European Union’s Horizon Europe Framework Programme for Research and Innovation, under the PATH project (Grant Agreement No. 101253520). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union, which cannot be held responsible for them.
@@ Appendix B Derivation of Glorot-inspired Initialization Scheme
In this appendix, we derive Eqs. (5) from the main text. Unlike the LeCun-inspired scheme, which focuses solely on variance preservation in the forward pass, the Glorot principle [Glorot and Bengio, 2010] requires that the variance of both activations and backpropagated gradients remain constant across layers. For analytical tractability, and following the standard assumption in this setting, we further approximate these constant values by unity,
@@
Equations (13) and (17) are the forward- and backward-pass constraints, respectively. Balancing them in the Glorot manner (i.e., by harmonic averaging) yields the standard deviations in Eq. (5) of the main text. As a sanity check, consider an MLP: the residual term is absent, and the linear layer followed by a nonlinearity can be viewed as a single basis function. For the common hyperbolic tangent activation, $\mu_{B}^{(0)}\approx\mu_{B}^{(1)}\approx 1$ [Glorot and Bengio, 2010], so our scheme reduces to
@@ C.1 Function Fitting
The KAN models used to fit these functions utilize spline basis functions of order $k=3$, defined over an augmented, uniform grid within the $\left[-1,1\right]$ domain [Liu et al., 2025]. Training is performed using the Adam optimizer with a fixed learning rate of $10^{-3}$, with the objective of minimizing the mean squared error between the predicted and reference function values. For each target function $f_{i}\left(x,y\right)$, with $i=1,\dots,5$, we generate 4,000 random input samples uniformly distributed over the domain $\left[-1,1\right]\times\left[-1,1\right]$, and calculate the corresponding outputs to serve as ground truth during training. To compute the relative $L^{2}$ error between the model predictions and reference solutions, we evaluate all trained models on a uniform $200\times 200$ grid covering the same domain.
@@ C.2 Forward PDE Problems
Since the Allen–Cahn equation has no analytic closed-form solution, we use the reference solution used in Wang et al. [2024], which is depicted in the left plot of Fig. 6.
@@
Similar to the Allen–Cahn equation, Burger’s equation has no analytic closed-form solution, therefore we use the reference solution used in Wang et al. [2024], which is depicted in the middle plot of Fig. 6.
@@
The PDE problems are solved using the Residual-Based Attention (RBA) weighting scheme [Anagnostopoulos et al., 2024] within the PIML framework [Raissi et al., 2019], where the training objective is defined as a sum of weighted residuals associated with the PDE differential operator and the boundary/initial condition operators. Specifically, we minimize
$\mathcal{L}\left(\boldsymbol{\theta}\right)=\frac{1}{N_{\text{pde}}}\sum_{i=1}^{N_{\text{pde}}}\left|\alpha^{(\mathrm{pde})}_{i}\,r^{(\mathrm{pde})}_{i}\left(\boldsymbol{\theta}\right)\right|^{2}+\frac{1}{N_{\text{bc}}}\sum_{i=1}^{N_{\text{bc}}}\left|\alpha^{(\mathrm{bc})}_{i}\,r^{(\mathrm{bc})}_{i}\left(\boldsymbol{\theta}\right)\right|^{2},$ (31)
@@ C.3 Feynman Dataset
As a third benchmark, we consider the subset of the Feynman dataset used in Section 4.3. The implementation details are identical to those of the function fitting benchmarks in Appendix C.1, with the exception of sampling. In this case, we generate 4,000 random input samples uniformly distributed over the domain $\left(-1,0\right)\cup\left(0,1\right)$, explicitly excluding the points 1-1, 0, and $1$ to avoid singularities in certain formulas.
@@ Appendix E Training Curves with Learning Rate Scheduling
In Section 4.2 of the main text, the training curves shown in Figures 1 and 2 were obtained using a fixed learning rate in order to isolate the effect of initialization, as initialization and learning-rate adaptability are known to interact (e.g., [Yang et al., 2021]). However, the fixed learning rate induces oscillations in the loss curves, particularly for the Glorot and power-law schemes in the larger architectures. To verify that these oscillations are purely an artifact of the constant learning rate, we repeat the same training experiments using a learning-rate scheduler.
@@ Appendix F Neural Tangent Kernel Analysis
In this work, we use NTK analysis [Jacot et al., 2018] to better understand the effect of initialization schemes on function fitting and PDE benchmarks, both in terms of stability and conditioning.
@@ F.1 NTK for PIML with RBA Weights
In this subsection, we derive the NTK formalism used in our PDE experiments. Specifically, we extend the standard NTK framework for PIML [Wang et al., 2022] to cover the RBA-weighted loss function of Eq. (31).
We denote the PDE and boundary/initial condition residuals at the $i$-th collocation point by $r^{(\mathrm{pde})}_{i}$ and $r^{(\mathrm{bc})}_{i}$, respectively, as in Appendix C.2. We may re-weight the loss function of Eq. (31) to follow Wang et al. [2022] and subsequently write it in vector form as
$\mathcal{L}(\boldsymbol{\theta})\;=\;\frac{1}{2}\,\big\|\tilde{\mathbf{r}}^{(\mathrm{pde})}(\boldsymbol{\theta})\big\|_{2}^{2}\;+\;\frac{1}{2}\,\big\|\tilde{\mathbf{r}}^{(\mathrm{bc})}(\boldsymbol{\theta})\big\|_{2}^{2},\qquad\tilde{\mathbf{r}}^{(\xi)}=\mathbf{A}^{(\xi)}\mathbf{r}^{(\xi)},$ (33)
where $\mathbf{r}^{(\xi)}$ stacks the residuals of type $\xi\in\{\mathrm{pde},\mathrm{bc}\}$, $\boldsymbol{\alpha}^{(\xi)}=(\alpha^{(\xi)}_{1},\dots,\alpha^{(\xi)}_{N_{\xi}})^{\top}$ are the RBA weights and $\mathbf{A}^{(\xi)}=\mathrm{diag}(\boldsymbol{\alpha}^{(\xi)})$. Throughout a single gradient step we treat $\boldsymbol{\alpha}^{(\xi)}$ as constants, as they are updated only between steps by Eq. (32), outside of the gradient descent scheme.
Let $\mathbf{J}^{(\xi)}(\boldsymbol{\theta})\in\mathbb{R}^{N_{\xi}\times P}$ be the Jacobian of the residuals with respect to the parameters, i.e., its $i$-th row is $\mathbf{J}^{(\xi)}_{i}(\boldsymbol{\theta})=\partial r^{(\xi)}_{i}(\boldsymbol{\theta})/\partial\boldsymbol{\theta}^{\top}$. For a parameter update $\Delta\boldsymbol{\theta}=-\eta\,\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})$, a first-order expansion around $\boldsymbol{\theta}$ yields,
$\Delta\tilde{\mathbf{r}}^{(\xi)}(\boldsymbol{\theta})\;=\;\mathbf{A}^{(\xi)}\Delta\mathbf{r}^{(\xi)}(\boldsymbol{\theta})\;\approx\;\mathbf{A}^{(\xi)}\mathbf{J}^{(\xi)}(\boldsymbol{\theta})\,\Delta\boldsymbol{\theta}.$ (34)
@@
$\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})\;=\;\sum_{i=1}^{N_{\mathrm{pde}}}\tilde{r}^{(\mathrm{pde})}_{i}(\boldsymbol{\theta})\,\nabla_{\boldsymbol{\theta}}\,\tilde{r}^{(\mathrm{pde})}_{i}(\boldsymbol{\theta})\;+\;\sum_{i=1}^{N_{\mathrm{bc}}}\tilde{r}^{(\mathrm{bc})}_{i}(\boldsymbol{\theta})\,\nabla_{\boldsymbol{\theta}}\,\tilde{r}^{(\mathrm{bc})}_{i}(\boldsymbol{\theta}).$ (35)
Since $\tilde{r}^{(\xi)}_{i}=\alpha^{(\xi)}_{i}\,r^{(\xi)}_{i}$ and $\boldsymbol{\alpha}^{(\xi)}$ is held fixed within the step,
$\nabla_{\boldsymbol{\theta}}\,\tilde{r}^{(\xi)}_{i}(\boldsymbol{\theta})\;=\;\alpha^{(\xi)}_{i}\nabla_{\boldsymbol{\theta}}\,r^{(\xi)}_{i}(\boldsymbol{\theta})\;=\;\alpha^{(\xi)}_{i}\big(\mathbf{J}^{(\xi)}_{i}(\boldsymbol{\theta})\big)^{\top}.$ (36)
Substituting $\Delta\boldsymbol{\theta}=-\eta\,\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})$ into Eq. (34) and grouping terms gives the linear dynamics
$\Delta\tilde{\mathbf{r}}^{(\xi)}(\boldsymbol{\theta})\;\approx\;-\,\eta\Bigg[$ $\underbrace{\big(\mathbf{A}^{(\xi)}\mathbf{J}^{(\xi)}(\boldsymbol{\theta})\big)\big(\mathbf{A}^{(\mathrm{pde})}\mathbf{J}^{(\mathrm{pde})}(\boldsymbol{\theta})\big)^{\top}}_{\displaystyle\widetilde{\mathbf{K}}^{(\xi,\mathrm{pde})}}\tilde{\mathbf{r}}^{(\mathrm{pde})}(\boldsymbol{\theta})$
$\quad+\;\underbrace{\big(\mathbf{A}^{(\xi)}\mathbf{J}^{(\xi)}(\boldsymbol{\theta})\big)\big(\mathbf{A}^{(\mathrm{bc})}\mathbf{J}^{(\mathrm{bc})}(\boldsymbol{\theta})\big)^{\top}}_{\displaystyle\widetilde{\mathbf{K}}^{(\xi,\mathrm{bc})}}\tilde{\mathbf{r}}^{(\mathrm{bc})}(\boldsymbol{\theta})\Bigg].$ (37)

### v3

@@ Abstract
Kolmogorov–Arnold Networks (KANs) are a recently introduced neural architecture that replace fixed nonlinearities with trainable activation functions, offering enhanced flexibility and interpretability. While KANs have been applied successfully across scientific and machine learning tasks, their initialization strategies remain largely unexplored. In this work, we study initialization schemes for spline-based KANs, proposing two theory-driven approaches inspired by LeCun and Glorot, as well as an empirical power-law family with tunable exponents. Our evaluation combines large-scale grid searches on function fitting and forward PDE benchmarks, an analysis of training dynamics through the lens of the Neural Tangent Kernel, and evaluations on a subset of the Feynman dataset. Our findings indicate that the Glorot-inspired initialization significantly outperforms the baseline in parameter-rich models, while power-law initialization achieves the strongest performance overall, both across tasks and for architectures of varying size. This work underscores initialization as a key factor in KAN performance and introduces practical strategies to improve it.
@@ Acknowledgments
S. R. and G. A. are supported by the Innovative Health Initiative Joint Undertaking (IHI JU) under grant agreement No 101253520. The JU receives support from the European Union’s Horizon Europe research and innovation programme and COCIR, EFPIA, Europa Bío, MedTech Europe, and Vaccines Europe.
LLM Usage
Large Language Models (LLMs) were used during peer review for grammar and syntax refinement only; all ideas, technical content, analyses and conclusions remain the authors’ work.
@@ Appendix B Derivation of Glorot-inspired Initialization Scheme
In this appendix, we derive Eqs. (5) from the main text. Unlike the LeCun-inspired scheme, which focuses solely on variance preservation in the forward pass, the Glorot principle (Glorot and Bengio, 2010) requires that the variance of both activations and backpropagated gradients remain constant across layers. For analytical tractability, and following the standard assumption in this setting, we further approximate these constant values by unity,
@@
Equations (13) and (17) are the forward- and backward-pass constraints, respectively. Balancing them in the Glorot manner (i.e., by harmonic averaging) yields the standard deviations in Eq. (5) of the main text. As a sanity check, consider an MLP: the residual term is absent, and the linear layer followed by a nonlinearity can be viewed as a single basis function. For the common hyperbolic tangent activation, $\mu_{B}^{(0)}\approx\mu_{B}^{(1)}\approx 1$ (Glorot and Bengio, 2010), so our scheme reduces to
@@ C.1 Function Fitting
The KAN models used to fit these functions utilize spline basis functions of order $k=3$, defined over an augmented, uniform grid within the $\left[-1,1\right]$ domain (Liu et al., 2025). Training is performed using the Adam optimizer with a fixed learning rate of $10^{-3}$, with the objective of minimizing the mean squared error between the predicted and reference function values. For each target function $f_{i}\left(x,y\right)$, with $i=1,\dots,5$, we generate 4,000 random input samples uniformly distributed over the domain $\left[-1,1\right]\times\left[-1,1\right]$, and calculate the corresponding outputs to serve as ground truth during training. To compute the relative $L^{2}$ error between the model predictions and reference solutions, we evaluate all trained models on a uniform $200\times 200$ grid covering the same domain.
@@ C.2 Forward PDE Problems
Since the Allen–Cahn equation has no analytic closed-form solution, we use the reference solution used in Wang et al. (2024), which is depicted in the left plot of Fig. 6.
@@
Similar to the Allen–Cahn equation, Burger’s equation has no analytic closed-form solution, therefore we use the reference solution used in Wang et al. (2024), which is depicted in the middle plot of Fig. 6.
@@
The PDE problems are solved using the Residual-Based Attention (RBA) weighting scheme (Anagnostopoulos et al., 2024) within the PIML framework (Raissi et al., 2019), where the training objective is defined as a sum of weighted residuals associated with the PDE differential operator and the boundary/initial condition operators. Specifically, we minimize
$\mathcal{L}\left(\bm{\theta}\right)=\frac{1}{N_{\text{pde}}}\sum_{i=1}^{N_{\text{pde}}}\left|\alpha^{(\mathrm{pde})}_{i}\,r^{(\mathrm{pde})}_{i}\left(\bm{\theta}\right)\right|^{2}+\frac{1}{N_{\text{bc}}}\sum_{i=1}^{N_{\text{bc}}}\left|\alpha^{(\mathrm{bc})}_{i}\,r^{(\mathrm{bc})}_{i}\left(\bm{\theta}\right)\right|^{2},$ (31)
@@ C.3 Feynman Dataset
As a third benchmark, we consider the subset of the Feynman dataset used in Section 4.3. The implementation details are identical to those of the function fitting benchmarks in Appendix C.1, with the exception of sampling. In this case, we generate 4,000 random input samples uniformly distributed over the domain $\left(-1,0\right)\cup\left(0,1\right)$, explicitly excluding the points 1-1, 0, and 11 to avoid singularities in certain formulas.
@@ Appendix E Training Curves with Learning Rate Scheduling
In Section 4.2 of the main text, the training curves shown in Figures 1 and 2 were obtained using a fixed learning rate in order to isolate the effect of initialization, as initialization and learning-rate adaptability are known to interact (e.g., (Yang et al., 2021)). However, the fixed learning rate induces oscillations in the loss curves, particularly for the Glorot and power-law schemes in the larger architectures. To verify that these oscillations are purely an artifact of the constant learning rate, we repeat the same training experiments using a learning-rate scheduler.
@@ Appendix F Neural Tangent Kernel Analysis
In this work, we use NTK analysis (Jacot et al., 2018) to better understand the effect of initialization schemes on function fitting and PDE benchmarks, both in terms of stability and conditioning.
@@ F.1 NTK for PIML with RBA Weights
In this subsection, we derive the NTK formalism used in our PDE experiments. Specifically, we extend the standard NTK framework for PIML (Wang et al., 2022) to cover the RBA-weighted loss function of Eq. (31).
We denote the PDE and boundary/initial condition residuals at the $i$-th collocation point by $r^{(\mathrm{pde})}_{i}$ and $r^{(\mathrm{bc})}_{i}$, respectively, as in Appendix C.2. We may re-weight the loss function of Eq. (31) to follow Wang et al. (2022) and subsequently write it in vector form as
$\mathcal{L}(\bm{\theta})\;=\;\frac{1}{2}\,\big\|\tilde{\mathbf{r}}^{(\mathrm{pde})}(\bm{\theta})\big\|_{2}^{2}\;+\;\frac{1}{2}\,\big\|\tilde{\mathbf{r}}^{(\mathrm{bc})}(\bm{\theta})\big\|_{2}^{2},\qquad\tilde{\mathbf{r}}^{(\xi)}=\mathbf{A}^{(\xi)}\mathbf{r}^{(\xi)},$ (33)
where $\mathbf{r}^{(\xi)}$ stacks the residuals of type $\xi\in\{\mathrm{pde},\mathrm{bc}\}$, $\bm{\alpha}^{(\xi)}=(\alpha^{(\xi)}_{1},\dots,\alpha^{(\xi)}_{N_{\xi}})^{\top}$ are the RBA weights and $\mathbf{A}^{(\xi)}=\mathrm{diag}(\bm{\alpha}^{(\xi)})$. Throughout a single gradient step we treat $\bm{\alpha}^{(\xi)}$ as constants, as they are updated only between steps by Eq. (32), outside of the gradient descent scheme.
Let $\mathbf{J}^{(\xi)}(\bm{\theta})\in\mathbb{R}^{N_{\xi}\times P}$ be the Jacobian of the residuals with respect to the parameters, i.e., its $i$-th row is $\mathbf{J}^{(\xi)}_{i}(\bm{\theta})=\partial r^{(\xi)}_{i}(\bm{\theta})/\partial\bm{\theta}^{\top}$. For a parameter update $\Delta\bm{\theta}=-\eta\,\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})$, a first-order expansion around $\bm{\theta}$ yields,
$\Delta\tilde{\mathbf{r}}^{(\xi)}(\bm{\theta})\;=\;\mathbf{A}^{(\xi)}\Delta\mathbf{r}^{(\xi)}(\bm{\theta})\;\approx\;\mathbf{A}^{(\xi)}\mathbf{J}^{(\xi)}(\bm{\theta})\,\Delta\bm{\theta}.$ (34)
@@
$\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})\;=\;\sum_{i=1}^{N_{\mathrm{pde}}}\tilde{r}^{(\mathrm{pde})}_{i}(\bm{\theta})\,\nabla_{\bm{\theta}}\,\tilde{r}^{(\mathrm{pde})}_{i}(\bm{\theta})\;+\;\sum_{i=1}^{N_{\mathrm{bc}}}\tilde{r}^{(\mathrm{bc})}_{i}(\bm{\theta})\,\nabla_{\bm{\theta}}\,\tilde{r}^{(\mathrm{bc})}_{i}(\bm{\theta}).$ (35)
Since $\tilde{r}^{(\xi)}_{i}=\alpha^{(\xi)}_{i}\,r^{(\xi)}_{i}$ and $\bm{\alpha}^{(\xi)}$ is held fixed within the step,
$\nabla_{\bm{\theta}}\,\tilde{r}^{(\xi)}_{i}(\bm{\theta})\;=\;\alpha^{(\xi)}_{i}\nabla_{\bm{\theta}}\,r^{(\xi)}_{i}(\bm{\theta})\;=\;\alpha^{(\xi)}_{i}\big(\mathbf{J}^{(\xi)}_{i}(\bm{\theta})\big)^{\top}.$ (36)
Substituting $\Delta\bm{\theta}=-\eta\,\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})$ into Eq. (34) and grouping terms gives the linear dynamics
$\Delta\tilde{\mathbf{r}}^{(\xi)}(\bm{\theta})\;\approx\;-\,\eta\Bigg[$ $\underbrace{\big(\mathbf{A}^{(\xi)}\mathbf{J}^{(\xi)}(\bm{\theta})\big)\big(\mathbf{A}^{(\mathrm{pde})}\mathbf{J}^{(\mathrm{pde})}(\bm{\theta})\big)^{\top}}_{\displaystyle\widetilde{\mathbf{K}}^{(\xi,\mathrm{pde})}}\tilde{\mathbf{r}}^{(\mathrm{pde})}(\bm{\theta})$
$\quad+\;\underbrace{\big(\mathbf{A}^{(\xi)}\mathbf{J}^{(\xi)}(\bm{\theta})\big)\big(\mathbf{A}^{(\mathrm{bc})}\mathbf{J}^{(\mathrm{bc})}(\bm{\theta})\big)^{\top}}_{\displaystyle\widetilde{\mathbf{K}}^{(\xi,\mathrm{bc})}}\tilde{\mathbf{r}}^{(\mathrm{bc})}(\bm{\theta})\Bigg].$ (37)